# Space to Policy: Scalable Brick Kiln Detection and Automatic Compliance Monitoring with Geospatial Data

### v1

@@ Abstract.
Air pollution is responsible for approximately 7 million deaths annually. The brick manufacturing industry significantly contributes to both economic development and environmental challenges, accounting for 8–14% of air pollution in India. Brick kilns provide employment to millions of workers, raising concerns related to the United Nations Sustainable Development Goals (UN SDGs), such as Goal 8.7 (modern slavery). Policymakers have implemented various regulations and compliance measures for brick kilns to mitigate air pollution and associated health effects. Emission inventories quantifying the contribution of different sectors, including but not limited to traffic, construction, brick kilns, and household burnings are critical for air quality modeling and source apportionment studies. However, the largely unorganized nature of the brick kiln sector necessitates labor-intensive survey efforts, hindering compliance assessment and emission inventory development. Recent efforts by air quality researchers have relied on manual annotation of brick kilns using satellite imagery to build emission inventories, but this approach lacks scalability. Machine-learning-based object detection methods have shown promise for detecting brick kilns; however, previous studies often rely on costly high-resolution imagery and fail to integrate the analysis with governmental policies. To address these limitations and support evidence-based decision-making, we developed a scalable machine-learning pipeline that detected and classified brick kilns across five states in the Indo-Gangetic Plain using a free-for-research, moderate-resolution satellite imagery. Our study covers an area of over 520,000 km2 with a population of 448 million. We validated our detections with on-ground surveys, achieving a high correlation of 0.94. Additionally, we performed automated compliance analysis based on government policies and found that a considerable number of brick kilns violate at least one compliance rule. We find that, in the Delhi airshed, stricter policy enforcement and heightened focus on air pollution have led to the adoption of more efficient brick kiln technologies. Using chemical transport modeling (WRF-CAMx), we conducted a PM2.5 source apportionment study, estimating the brick kiln contribution to be 8% in the Delhi airshed. Furthermore, over 30 million people live within 800 meters of brick kilns, contravening government policies. Acknowledging the socio-economic importance of the brick kiln sector, this study highlights the need for inclusive policies that balance environmental sustainability with the livelihoods of workers. By leveraging advances in machine learning and satellite imagery, this work provides tools to foster cleaner air, healthier communities, and a more sustainable future.
@@ 1. Introduction
Initially, many brick kilns operated with Fixed Chimney Bull’s Trench Kiln (FCBK), an older technology. Studies show that converting to Zigzag technology can reduce CO and particulate matter emissions by 70% (Tibrewal et al., 2023). FCBKs and Zigzag kilns account for 75% of the brick production in India (Rajarathnam et al., 2014). The National Clean Air Program (NCAP), launched in 2019, introduced guidelines to mitigate air pollution and identified 130 cities as “non-attainment cities” due to non-compliance with air quality guidelines. Indian central government and state governments have established compliance rules for brick kilns to regulate their operations (Bri, [n. d.]). Source apportionment studies and air quality modeling rely on accurate emission inventory (Guttikunda et al., 2019). Developing emission inventories and monitoring compliance require accurate geo-mapping of brick kilns and their technologies. However, brick kilns are small-scale, unorganized units, making geo-mapping through field surveys resource-intensive.
@@
This article is structured as follows: Section 2 provides an overview of brick kilns, their impact on air pollution and alignment with UN SDGs, relevant Indian government policies, and object detection models. Section 8 reviews related studies on object detection from aerial imagery, the application of computer vision for UN SDGs, and satellite-based brick kiln detection. In Section 3, we outline the big data resources used for automated compliance monitoring and population exposure estimation. Section 4 details the process of creating a geo-located brick kiln dataset using object detection on satellite imagery. Section 5 presents results from automated compliance monitoring. Section 6 explores the experiments linking air pollution and health effects associated with brick kilns. Section 7 discusses potential policy improvements, and finally, we address the study’s limitations, propose future work, and conclude.
Reproducibility Statement: our work is fully reproducible and we will release all the code and resource links to reproduce the results in this work upon acceptance.
@@ 2.4. Government Policies
The Environment (Protection) Amendment Rules, 2022 (Bri, [n. d.]), mandated the conversion of all brick kilns within 10 km of non-attainment cities to cleaner technologies, including Zigzag technology, vertical shaft kilns, or piped natural gas, by 2024. The rules also specified:
@@ 3.1.1. Planet Labs
We utilize satellite imagery from Planet Labs in this work (Labs, [n. d.]). Planet offers multiple satellite products, including 3-meter resolution daily imagery, 4.77-meter monthly and quarterly mosaics111https://desktop.arcgis.com/en/arcmap/latest/manage-data/raster-and-images/what-is-a-mosaic.htm. Imagery is provided through Planet’s Research and Education Program (Labs, [n. d.]) and can be freely downloaded via the Mosaics API (Pla, [n. d.]). Our study area spans five Indian states, covering over 520,000 km² (15.8% of India) area. The large sized images are divided into overlapping sized crops with a 64-pixel overlap (300 meters), ensuring that brick kilns do not get cut at the image edges. We selected imagery from the first quarter of 2024 to align with the peak operational season and favorable weather conditions for cloud-free imagery. Although Planet imagery offers better resolution than Sentinel-2, it remains insufficient for identifying kiln technologies. To overcome this, we annotated high-resolution imagery and transferred the labels to Planet imagery, as detailed in the next section.
@@ 3.2. OpenStreetMap
OpenStreetMap (OSM) (OpenStreetMap contributors, 2024a), a community-driven initiative, offers a vast array of geographic datasets contributed by volunteers globally. These datasets include detailed information on infrastructure and natural features, enabling extensive spatial analysis. Using the Geofabrik platform (OpenStreetMap contributors, 2024b), we extracted geometry data for features such as railway tracks, highways, rivers, schools, and habitation. We utilize this data for distance-based compliance monitoring. Table 1 summarizes the extracted features and provides the filters we use to extract the data.
@@ 3.4. Population Data
We used the LandScan global population dataset (Lebakula et al., 2024), produced by Oak Ridge National Laboratory, at a resolution of 0.0083 degrees. We download the 2023 version of the dataset. We utilize this dataset to analyze population exposure due to brick kilns’ air pollution, as detailed in Section 6.3.
@@ 4.1. Initial Data Preparation
Across these four regions, we manually scanned over 15,018 km2 area and identified 1,621 brick kilns, completing the annotations in 125 hours. Table 2 provides region-wise details of area, annotation time, and class-wise distribution. The selected regions are visualized in Figure 7 in the Appendix.
@@ 4.2. Annotation Process
The annotation process entails labeling brick kilns with oriented bounding boxes (OBBs) and assigning a category to represent the kiln’s technology. We developed a custom labeling interface using Leafmap (Wu, 2021), as shown in Figure 2, utilizing the Esri Wayback Satellite Imagery WMS layer. The region of interest was divided into a 1 km2 grid (Figure 2(a)), and each grid cell was manually inspected. Within each cell, we draw OBBs around the identified kilns (Figure 2(b)). After completing the annotations, the labeled data were overlaid onto cropped, geo-referenced Planet images. Since both Esri and Planet imagery are in the Web Mercator projection (EPSG:3857), the OBBs remained undistorted during reprojection from Esri to Planet imagery, as demonstrated in Figure 1.
@@ 4.3. Model Selection
We utilize YOLO-OBB models provided by Ultralytics (Jocher et al., 2023) for our experiments. These models deliver performance comparable to state-of-the-art OBB detection frameworks. YOLO’s algorithm enables faster inference by detecting objects in a single forward pass, making it well-suited for scaling predictions across large geographic areas. We evaluate all OBB variants (‘YOLOv8’ and ‘YOLO11’) across different sizes (‘n’: nano, ‘s’: small, ‘m’: medium, ‘l’: large, ‘x’: extra-large). The dataset described in Table 2 is combined and split into stratified subsets with an 80:10:10 ratio for training, validation, and testing. Stratification ensures a proportional representation of all classes in each subset. The training set (‘train’) is used to train the models, while the validation set (‘validation’) guides performance monitoring and early stopping. The test set (‘test’) evaluates final model performance. We measure model accuracy using the ‘Weighted mAP50’ (mean Average Precision at 50% Intersection over Union) metric, which adjusts the standard mAP50 by weighting class-specific average precisions (APs) inversely to their test dataset counts. Table 3 summarizes the performance of various models on the test dataset. Among the evaluated models, yolo11m-obb achieves the highest performance and is therefore used for brick kiln detection in subsequent experiments.
@@ 4.4. Iterative Dataset Building Process
in CFCBK_106.png, CFCBK_12.png, CFCBK_131.png, CFCBK_39.png, CFCBK_5.png, CFCBK_50.png, FCBK_0.png, FCBK_1.png, FCBK_10.png, FCBK_13.png, FCBK_14.png, FCBK_17.png, Zigzag_11.png, Zigzag_16.png, Zigzag_18.png, Zigzag_2.png, Zigzag_20.png, Zigzag_22.png
@@ 5.1. Distance-based Compliance Monitoring
The Environment (Protection) Amendment Rules, 2022 (Bri, [n. d.]), outline compliance requirements for brick kiln operations across all states of India. Among these, two key distance-based siting rules are: i) brick kilns must maintain a minimum distance of 800 meters from habitation and fruit orchards; ii) brick kilns must be at least 1 kilometer apart from one another. States are permitted to establish additional or stricter siting criteria. For our analysis, if both national and state-specific criteria exist, we prioritize state-specific rules. We examined the compliance criteria for Uttar Pradesh, Bihar, West Bengal, Haryana, and Punjab. Using the big data resources described in Section 3, we automatically computed distances between points of interest (e.g., habitation, hospitals, other kilns) and brick kilns. A summary of the results is provided in Table 5, with the corresponding distance thresholds detailed in Table 9 in the Appendix. We visualize the locations of violating brick kilns in Figure 8. Based on our analysis, the following patterns emerge:

### v3

@@ Abstract.
Air pollution kills 7 million people annually. The brick kiln sector significantly contributes to economic development but also accounts for 8-14% of air pollution in India. Policymakers have implemented compliance measures to regulate brick kilns. Emission inventories are critical for air quality modeling and source apportionment studies. However, the largely unorganized nature of the brick kiln sector necessitates labor-intensive survey efforts for monitoring. Recent efforts by air quality researchers have relied on manual annotation of brick kilns using satellite imagery to build emission inventories, but this approach lacks scalability. Machine-learning-based object detection methods have shown promise for detecting brick kilns; however, previous studies often rely on costly high-resolution imagery and fail to integrate with governmental policies. In this work, we developed a scalable machine-learning pipeline that detected and classified 30638 brick kilns across five states in the Indo-Gangetic Plain using free, moderate-resolution satellite imagery from Planet Labs. Our detections have a high correlation with on-ground surveys. We performed automated compliance analysis based on government policies. In the Delhi airshed, stricter policy enforcement has led to the adoption of efficient brick kiln technologies. This study highlights the need for inclusive policies that balance environmental sustainability with the livelihoods of workers.
@@ 1. Introduction
Initially, many brick kilns operated with Fixed Chimney Bull’s Trench Kiln (FCBK), an older technology. Studies show that converting to Zigzag technology can reduce CO and particulate matter emissions by 70% (Tibrewal et al., 2023). FCBKs and Zigzag kilns account for 75% of the brick production in India (Rajarathnam et al., 2014). The National Clean Air Program (NCAP), launched in 2019, introduced guidelines to mitigate air pollution and identified 130 cities as “non-attainment cities” due to non-compliance with air quality guidelines. Indian central government and state governments have established compliance rules for brick kilns to regulate their operations (Bri, [n. d.]). Source apportionment studies and air quality modeling rely on accurate emission inventory (Guttikunda et al., 2019b). Developing emission inventories and monitoring compliance require accurate geo-mapping of brick kilns and their technologies. However, brick kilns are small-scale, unorganized units, making geo-mapping through field surveys resource-intensive.
@@
This article is structured as follows: Section 2 provides an overview of brick kilns, their impact on air pollution, alignment with UN SDGs, relevant Indian government policies, and object detection models. Section 8 reviews related studies on object detection from aerial imagery, the application of computer vision for UN SDGs, and satellite-based brick kiln detection. In Section 3, we outline the big data resources used for automated compliance monitoring and population exposure estimation. Section 4 details the process of creating a geo-located brick kiln dataset using object detection on satellite imagery. Section 5 presents results from automated compliance monitoring. Section 6 explores the experiments linking air pollution and health effects associated with brick kilns. Section 7 discusses socio-economic, data and policy implications, and finally, we discuss our study’s limitations, propose future work, and provide a conclusion.
Reproducibility Statement: Code and resources to reproduce our work can be found at our project page111https://sustainability-lab.github.io/brick-kilns/
@@ 2.4. Government Policies
The Environment (Protection) Amendment Rules, 2022 (Bri, [n. d.]), mandated the conversion of all brick kilns within 10 km of non-attainment cities to cleaner technologies, including Zigzag technology, vertical shaft kilns, or piped natural gas, by 2024.
2.5. Brick Kiln Siting Rules
Industrial activities can significantly impact the environment through cumulative emissions, raw material extraction, and industrial waste. Therefore, industries must be sited with a balance between economic and environmental considerations. Several India states have established brick kiln siting regulations. One such example is the Brick Kiln Siting Rules, 2012 of Uttar Pradesh (Uttar Pradesh Government, 2012). Air pollution from brick kilns, particularly the presence of sulfur dioxide (SO2) and hydrogen fluoride (HF), negatively impacts the flowering of fruit trees (e.g., mangoes), leading to reduced fruit yield (Ahmad et al., 2012). To mitigate this, the rules specify a minimum distance from fruit orchards. To reduce the impact of air pollution on human health, the rules also mandate a minimum distance from residential areas, hospitals, and schools. The primary source of brick earth, the raw material for brick-making, comes from mining topsoil from agricultural fields surrounding the kiln, typically up to a 2 meters deep. This practice can alter natural rainwater flow and lead to land erosion. To address this, the kiln siting rules specify minimum distances from railway tracks, state highways, and national highways. To prevent the formation of large brick kiln clusters, which can cause a significant cumulative environmental impacts, the Uttar Pradesh kiln siting rules establish a minimum distance between adjacent kilns. The following distance-based criteria have been established by the central government for all states (Bri, [n. d.]):
@@ 3.1.1. Planet Labs
We utilize satellite imagery from Planet Labs in this work (Labs, [n. d.]). Planet offers multiple satellite products, including 3-meter resolution daily imagery, 4.77-meter monthly and quarterly mosaics222https://desktop.arcgis.com/en/arcmap/latest/manage-data/raster-and-images/what-is-a-mosaic.htm. Imagery is provided through Planet’s Research and Education Program (Labs, [n. d.]) and can be freely downloaded via the Mosaics API (Pla, [n. d.]). Our study area spans five Indian states, covering over 520,000 km² (15.8% of India) area. The large sized images are divided into overlapping sized crops with a 64-pixel overlap (300 meters), ensuring that brick kilns do not get cut at the image edges. We selected imagery from the first quarter of 2024 to align with the peak operational season and favorable weather conditions for cloud-free imagery. Although Planet imagery offers better resolution than Sentinel-2, it remains insufficient for identifying kiln technologies. To overcome this, we annotated high-resolution imagery and transferred the labels to Planet imagery, as detailed in the next section.
@@ 3.2. OpenStreetMap
OpenStreetMap (OSM) (OpenStreetMap contributors, 2024a), a collaborative, community-driven initiative, provides a comprehensive repository of geographic datasets contributed by volunteers worldwide. These datasets contain detailed information on infrastructure and natural features, facilitating extensive spatial analysis. Using the Geofabrik platform (OpenStreetMap contributors, 2024b), we extracted geometry data for features such as railway tracks, highways, rivers, schools, and habitation. This data is utilized for distance-based compliance monitoring. Table 1 provides a detailed summary of the extracted features and the filters applied during the extraction process. The data is obtained by downloading zip files provided by the Geofabrik platform and extracting the data based on specific filters. Researchers can replicate our data collection process using the filters outlined in Table 1, ensuring transparency and reproducibility in our methodology.
@@ 3.4. Population Data
We used the LandScan global population dataset (Lebakula et al., 2024), produced by Oak Ridge National Laboratory, at a resolution of 0.0083 degrees. This dataset is updated once every year. We downloaded the 2023 version of the dataset. We used this dataset to analyze population exposure due to brick kilns’ air pollution, as detailed in Section 6.3. To determine the population within km of brick kilns, we summed all points falling within km radius of a brick kiln and ensured each point was counted only once to avoid duplication. This approach allows for accurate estimation of population exposure while maintaining data integrity.
@@ 4. Brick Kiln Dataset
In this section, we elaborate on the process of initial data preparation, model selection, brick kiln detection and external validation with surveys.
@@ 4.1. Initial Data Preparation
As shown in Table 2, we manually scanned over 15,018 km2 area and identified 1,621 brick kilns across the four regions, completing the annotations in 125 hours. Table 2 provides region-wise details of area, annotation time, and class-wise distribution. The selected regions are illustrated in Figure 2.
@@ 4.2. Annotation Process
The annotation process involves labeling brick kilns with oriented bounding boxes (OBBs) and assigning a category to represent the kiln’s technology. We developed a custom labeling interface using the Leafmap library (Wu, 2021), as shown in Figure 3, which utilizes the Esri Wayback Satellite Imagery WMS layer. The region of interest was divided into a set of grids (Figure 3(a)), and each grid cell was manually inspected for labeling. If a grid cell is too small, the annotation process becomes excessively time-consuming, whereas a grid cell too large may lead to missed kilns during manual inspection. A grid size of 1 km2 was found to provide a good tradeoff between speed and accuracy of annotation. Within each cell, we draw OBBs around the identified kilns (Figure 3(b)). The annotations created via our interface are geo-referenced, meaning that the coordinates of the bounding boxes are geo-coordinates and can thus be overlaid on any geo-referenced imagery. After completing the annotations, the labeled data were overlaid onto cropped, geo-referenced Planet images. Both Esri and Planet imagery are in the Web Mercator projection (EPSG:3857), and thus the OBBs remain undistorted after the reprojection from Esri to Planet imagery, as demonstrated in Figure 1.
@@ 4.3. Model Selection
We utilize YOLO-OBB models implemented by Ultralytics (Jocher et al., 2023) for our experiments. These models deliver performance comparable to state-of-the-art OBB detection frameworks. YOLO’s algorithm enables faster inference by detecting objects in a single forward pass, making it well-suited for scaling predictions across large geographic areas. We evaluate all OBB variants (‘YOLOv8’ and ‘YOLO11’) across different sizes (‘n’: nano, ‘s’: small, ‘m’: medium, ‘l’: large, ‘x’: extra-large). The dataset described in Table 2 is combined and split into stratified subsets with an 80:10:10 ratio for training, validation, and testing. Stratification ensures a proportional representation of all classes in each subset. The training set (‘train’) is used to train the models, while the validation set (‘validation’) guides performance monitoring and early stopping. The test set (‘test’) evaluates final model performance. We measure model accuracy using the ‘Weighted mAP50’ (mean Average Precision at 50% Intersection over Union) metric, which adjusts the standard mAP50 by weighting class-specific average precisions (APs) inversely to their test dataset counts. Table 3 summarizes the performance of various models on the test dataset. Among the evaluated models, yolo11m-obb achieves the highest performance and is therefore used for brick kiln detection in subsequent experiments.
@@
4.4. Out of the Region Performance
Deep learning models are not guaranteed to perform well when the test data differs significantly from the training data distribution. To evaluate the model’s robustness and estimate exclusion errors (i.e., undetected kilns), we conducted a ‘Leave One Region Out’ experiment. This involved using four initially labeled regions (described in Section 4.1) and iteratively designating one region for testing while training the best-performing model (as identified in Table 3) on the remaining regions. The results, presented in Table 4, reveal that the model’s mAP scores significantly drop in this setup. The Precision and Recall columns represent inclusion and exclusion errors, respectively (higher is better). Notably, Recall is consistently lower than Precision, indicating that exclusion errors are more prevalent than inclusion errors. This underscores the limited robustness of detection models when applied to regions distant from the training data. Although addressing this challenge is beyond the current paper’s scope, we mitigated it during the data-building process by iteratively incorporating labeled data from all regions to enhance model robustness. These findings highlight the importance of diverse and representative training data for improving model generalization. Future work could explore advanced techniques, such as domain adaptation to further address this limitation.
| Test Region | CFCBK | FCBK | Zigzag | Weighted mAP | TP | FP | FN | P | R |
| Delhi Airshed | 0.00 | 0.11 | 0.32 | 0.04 | 317 | 421 | 632 | 0.43 | 0.33 |
| Lucknow Airshed | 0.32 | 0.01 | 0.54 | 0.32 | 221 | 206 | 275 | 0.52 | 0.45 |
| West Bengal Small | 0.00 | 0.02 | 0.28 | 0.16 | 64 | 83 | 142 | 0.44 | 0.31 |
| Ahmedabad 10 km buffer | 0.26 | 0.06 | 0.00 | 0.18 | 18 | 47 | 131 | 0.28 | 0.12 |
@@ 4.6. External Validation
| District | Survey counts | Our counts |
| Baghpat | 300 | 432 |
| Bhiwani | 52 | 108 |
| Bulandshahr | 96 | 377 |
| Faridabad | 100 | 95 |
| GautamBudhNagar | 12 | 91 |
| Ghaziabad | 142 | 248 |
| Gurugram | 9 | 5 |
| Hapur | 104 | 127 |
| Jhajjar | 253 | 287 |
| Jind | 55 | 117 |
| Karnal | 89 | 85 |
| Mahendergarh | 37 | 57 |
| Meerut | 130 | 202 |
| Mewat | 62 | 77 |
| Muzaffarnagar | 105 | 341 |
| Palwal | 130 | 122 |
| Panipat | 68 | 74 |
| Rewari | 82 | 70 |
| Rohtak | 46 | 94 |
| Shamli | 86 | 153 |
| Sonipat | 194 | 204 |
| Total | 2152 | 3366 |
| District | Survey counts | Our counts |
| Bankura | 380 | 207 |
| Barddhaman | 396 | 488 |
| Howrah | 206 | 92 |
| Hugli | 208 | 126 |
| Maldah | 276 | 140 |
| Murshidabad | 600 | 467 |
| Nadia | 530 | 291 |
| North 24 Parganas | 387 | 370 |
| Pashchim Medinipur | 158 | 80 |
| Purba Medinipur | 317 | 244 |
| Puruliya | 168 | 89 |
| South 24 Parganas | 269 | 191 |
| Total | 3895 | 2785 |
| District | Survey counts | Our counts |
| Gopalganj | 264 | 201 |
| Nawada | 321 | 244 |
| Pashchim Champaran | 281 | 206 |
| Purba Champaran | 465 | 380 |
| Siwan | 349 | 229 |
| Total | 1680 | 1260 |
@@
CPCB submitted a compliance affidavit for Delhi-NCR brick kilns to the Supreme court of India in 2022333Diary No. 20331/2021, 18213/2021. We derived the counts of brick kilns in each district within Delhi-NCR and compared them against our detections in Table 8 and Figure 7. Figure 7(c) shows that our detections are higher compared to the counts derived from the affidavit. It indicates that between the survey time (around 2022) and our detection time (first quarter, 2024), many new brick kilns are established in the region. Similarly we derived the counts from a few districts of Bihar from the UNDP GeoAI report and reported the comparison with our counts in Figure 8 and Table 8. In another case involving West Bengal Pollution Control Board444Application No. 60/2020/EZ, a status report of conversion to Zigzag technology was submitted in the court. We derived the counts for the districts of West Bengal from the same UNDP report and presented the comparison with our counts in Figure 9 and Table 8. In Bihar and West Bengal, our counts are lower compared to the surveyed counts in most districts, which could either indicate that our exclusion errors are higher or number of kilns have reduced in the region. Overall, we observe a notable alignment of our detected counts with the reported counts in the surveys, affidavits and reports.
@@ 5.1. Distance-based Compliance Monitoring
The Environment (Protection) Amendment Rules, 2022 (Bri, [n. d.]), outline compliance requirements for brick kiln operations across all states of India. Among these, two key distance-based siting rules are: i) brick kilns must maintain a minimum distance of 800 meters from habitation and fruit orchards; ii) brick kilns must be at least 1 kilometer apart from one another. States are permitted to establish additional or stricter siting criteria. For our analysis, if both national and state-specific criteria exist, we prioritize state-specific rules.
We examined the compliance criteria for Uttar Pradesh (Uttar Pradesh Government, 2012), Bihar, West Bengal, Haryana, and Punjab. Using the big data resources described in Section 3, we automatically computed distances between points of interest (e.g., habitation, hospitals, other kilns) and brick kilns. A summary of the results is provided in Table 10, with the corresponding distance thresholds detailed in Table 9. We illustrate the locations of violating brick kilns in Figure 10. Based on our analysis, the following patterns emerge:
@@
| Criterion | Uttar Pradesh | Bihar | West Bengal | Haryana | Punjab |
| Inter kiln | 800 | 1000 | 300 | 1000 | 1000 |
| Habitation | 1000 | 800 | 800 | 800 | 500 |
| National highway | 300 | 300 | 200 | - | - |
| River | - | 500 | 200 | - | - |
| State highway | 300 | 200 | 200 | - | 100 |
| District highway | 100 | - | - | - | - |
| Railway | 200 | 200 | 200 | - | - |
| Nature reserve | 5000 | - | 5000 | 1000 | - |
| Orchard | 800 | 800 | 800 | 800 | 800 |
| Wetland | - | 500 | - | - | - |
| School | 1000 | 800 | 1000 | 1000 | - |
| Religious places | 1000 | - | 1000 | - | - |